OpenStudy (anonymous):

byron is 3 years older than doug. the product of their ages is 40. how old is doug?

OpenStudy (bibby):

Doug's age is some unknown x. Byron is 3 years older than that so he'd be x+3; multiply these 2 and set them equal to 40. Solve.

OpenStudy (anonymous):

Dougs age = x Byron = (x + 3) yrs Multiply the two ages to get 40: x * (x + 3) = 40 x^2 + 3x - 40 = 0 (x+8)(x-5) = 0 Factor: x = -8 or x = 5 The age cannot be negative so he is 5 yrs old